# On Utility Maximization under Multivariate Fake Stationary Affine Volterra Models

### 11 Mar 2026 v1

@@ 1 Introduction
The modeling of asset price dynamics has undergone a paradigm shift with the empirical observation that both implied and realized volatilities of major financial indices exhibit significantly rougher sample paths (Gatheral et al., 2018) than those generated by classical Brownian-motion-based models. This observation has sparked a rapidly expanding body of research on rough volatility models. Motivated by the widespread practical success of the celebrated Heston (1993) stochastic volatility framework, several rough extensions have been proposed, most notably the rough Heston model (El Euch and Rosenbaum, 2019) builds on market microstructure and its generalisation, the Volterra Heston model in Abi Jaber et al. (2019). Recent remarkable advances include the introduction of the so-called fake stationary Volterra Heston models in Gnabeyeu et al. (a, b) with the aim of providing a unified and consistent framework that captures both short- and long-maturity behaviors and enables robust fitting across the entire term structure. This broader class of volatility models, encompassing the aforementioned specifications, is obtained by modeling the volatility process as a stochastic Volterra equation of convolution type with a time-dependent diffusion coefficient. Therefore, this paper focuses on the financial market within the fake stationary affine Volterra model.
@@
Motivated by several important empirical stylized facts about real financial markets such as choice among multiple assets, rough volatility behavior, correlations across stocks or assets and leverage effects (i.e., correlation between a stock and its volatility), multivariate rough volatility models have recently been developed ; see, e.g., Abi Jaber et al. (2019); Tomas and Rosenbaum (2021). In Aichinger and Desmettre (2021), the authors analyze the Merton portfolio problems with power utility for a class of multivariate Volterra Heston models that features both inter-asset correlations and correlation between a stock and its volatility. In the present paper, we solve the Merton portfolio problem for investors with power and exponential utility functions, within the class of the so-called fake stationary multivariate affine Volterra models.
Main contributions. Building upon recent developments in volatility and Volterra models (Gnabeyeu et al., 2024, a, b) and motivated by recent works and advances on multivariate Volterra volatility modeling Abi Jaber et al. (2019); Tomas and Rosenbaum (2021); Aichinger and Desmettre (2021), the primary objective of this paper is to advance the literature on utility maximization along two main directions:
@@
Organization of the Work. The outline of the paper is as follows: Section 2 gives an overview of the model which is needed throughout the paper: We introduce the multi asset financial market, where volatility is modeled by a multivariate class of fake stationary Volterra square root process, and we state the optimization problem. For such a market model we consider in section 3 two different approaches to solve the Merton portfolio problem. More precisely, in section 3 we investigate the classical problem of maximizing the expected utility of terminal wealth in a multi-asset fake stationary Volterra–Heston volatility market, for both power and exponential utility preferences. We first adapt the martingale distortion transformation used in Fouque and Hu (2019); Han and Wong (2020) to the (degenerate) multivariate case. However, as it is pointed out in Abi Jaber et al. (2021); Aichinger and Desmettre (2021), this only works if the correlation structure is highly degenerate. Inspired by the techniques used in Bäuerle and Li (2013); Aichinger and Desmettre (2021), we then provide a solution for the Merton portfolio optimization problem for a more general correlation structure using a verification argument. In Section 4, we demonstrate the practical implications of our findings through numerical experiments based on a two-dimensional fake stationary rough Heston model. Finally, Section 5 is devoted to the proofs of the main results.
@@ 2 Preliminaries: Multivariate fake stationary affine Volterra models
$\|b(t,x)\|+\||\sigma(t,x)|\|\leq C_{b,\sigma}(1+\|x\|),\quad\text{for all }t\in[0,T]\text{ and }x\in\mathbb{R}.$
@@
Assume that $K$ is diagonal with scalar kernels $K_{i}$ on the diagonal for $i=1,\ldots,d$ that is completely monotone on $(0,\infty)$ and satisfy for any $T>0$,
@@
    Assume that the kernel $K_{i}$ is strictly positive and satisfy:
@@
Remark: For $i=1,\ldots,d$ , as $K_{i}$ is completely monotone on $(0,\infty)$ and not identically zero, we have that $K_{i}$ is nonnegative, not identically zero, non-increasing and continuous on $(0,\infty)$, and it follows from (B. Gripenberg and Saavalainen, 1990, Theorem 5.5.4) that $K_{i}$ has a resolvent of the first kind $r_{i}$ which is nonnegative and non-increasing in the sense that $s\mapsto r_{i}([s,s+t])$ is non-increasing for all $t\geq 0$.
In the case of $\alpha$- fractional kernel (corresponding to $K_{i}=K_{\alpha_{i}}$ with $\alpha_{i}\in[\frac{1}{2},1)$), by Gnabeyeu et al. (b) Equation (2.2) admits at least a unique-in-law positive weak solution as a scaling limit of a sequence of time-modulated Hawkes processes with heavy-tailed kernels in a nearly unstable regime. Moreover, under assumption 2.1 for some $p>0$, a solution $t\mapsto V_{t}$ to Equation (2.2) starting from $V_{0}$ has a $\big(\delta\wedge\vartheta\wedge\frac{\beta-1}{2\beta}-\eta\big)$-Hölder pathwise continuous modification on $\mathbb{R}_{+}^{d}$ for sufficiently small $\eta>0$ and satisfying (among other properties),
@@
This covers, for instance, constant non-negative kernels, fractional kernels of the form $\frac{t^{\alpha-1}}{\Gamma(\alpha)}\mathbf{1}_{\mathbb{R}_{+}}$ with $\alpha\in(\frac{1}{2},1]$, exponentially decaying kernels ${\rm e}^{-\beta t}$ with $\beta>0$ and more generally the gamma kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}e^{-\beta t}\mathbf{1}_{\mathbb{R}_{+}}$ with $\alpha\in\left(\tfrac{1}{2},1\right]$ and $\beta\geq 0$ ( see e.g. (Gnabeyeu and Pagès, 2025, Propositions 6.1 and 6.3)).
@@
$\ \mathbb{E}[V_{0}^{i}]=\frac{1-a_{i}}{1-a_{i}\varphi_{\infty}^{i}}\frac{\mu_{\infty}^{i}}{\lambda_{i}}:=x_{\infty}^{i}\qquad\text{and}\qquad\forall\,t\geq 0,\qquad\varphi^{i}(t)=1-\lambda_{i}\int_{0}^{t}K_{i}(t-s)\left(\frac{\mu^{i}(s)}{\lambda_{i}x_{\infty}^{i}}-1\right)\,\,\mathrm{d}s.$ (2.10)
@@
$\textit{($E_{\lambda_{i},c_{i}}$)}:\;\forall\,t\geq 0,\;c_{i}\lambda_{i}^{2}\big(1-(\varphi^{i}(t)-(f_{\lambda_{i}}*\varphi^{i})_{t})^{2}\big)=(f_{\lambda_{i}}^{2}*\varsigma^{i2})(t)\;\textit{where}\;c_{i}=\frac{v_{0}^{i}}{\nu_{i}^{2}x_{\infty}^{i}}\;\textit{i.e.}\;\varsigma^{i}=\varsigma^{i}_{\lambda_{i},c_{i}}.$ (2.12)
Proof : This is a straightforward extension to the multi-dimensional setting of (Gnabeyeu and Pagès, 2025, Proposition 3.4 and Theorem 3.5) (see also (Gnabeyeu et al., a, Proposition 4.2 and 4.4)).
@@
Set $E_{D,c}=\bigcup_{i=1}^{d}E_{\lambda_{i},c_{i}}$. From now on, we will assume that there exists a unique positive bounded Borel solution $\varsigma=\varsigma_{D,c}$ on $(0,+\infty)$ of the system of equation $(E_{D,c})$ so that, the corresponding time-inhomogeneous Volterra square root equation (2.2) is refered to as a Multivariate Stabilized Volterra Cox-Ingersoll-Ross (CIR) equation or as a Multivariate fake stationary Volterra CIR equation if, in addition, equation (2.10) holds.
@@ 2.2 Formulation of the stochastic Market model
We assume that $\sigma$ in (2.14) is given by $\sigma=\sqrt{\operatorname{diag}(V)}$, where the $\mathbb{R}^{d}_{+}$–valued scaled process $V$ is defined in (2.2) with $\varsigma=\varsigma_{D,c}$ and Equation (2.10) holds true. We will be chiefly interested in the case where $\lambda_{t}$ is linear in $\sigma_{t}$. More specifically, the the market price of risk (risk premium) is assumed to be in the form $\lambda$ == $\big(\theta_{1}\sqrt{V^{1}},\ldots,\theta_{d}\sqrt{V^{d}}\big)^{\top}$, for some constant $\theta_{i}\geq 0$, so that the dynamics for the stock prices (2.14) reads following Kraft (2005); Abi Jaber et al. (2019)
@@
We state the following existence and uniqueness result from Gnabeyeu et al. (b) which is extended to the multi-dimensional setting.
@@
((Gnabeyeu et al., b, Theorem 3.1 and Remark on Theorem 3.2)). Under Assumption 2.1, the stochastic Volterra equation (2.16)-(2.2) has a unique in law continuous $\mathbb{R}^{d}_{+}\times\mathbb{R}^{d}_{+}$-valued weak solution $(S,V)$ for any initial condition $(S_{0},V_{0})\in\mathbb{R}^{d}_{+}\times\mathbb{R}^{d}_{+}$ defined on some filtered probability space $(\Omega,\mathcal{F},(\mathcal{F})_{t\geq 0},\mathbb{P})$ such that
@@
From now on, we set $g_{0}(t):=\varphi(t)V_{0}+\int_{0}^{t}K(t-s)\mu(s)ds$ and $dZ_{t}=DV_{t}dt+\nu\sqrt{\operatorname{diag}(V_{t})}dW_{t},\;\forall t\geq 0$ so that Equation (2.2) reads
@@
This also provides an extension to the inhomogeneous setting of the models considered in Abi Jaber et al. (2019, 2021); Aichinger and Desmettre (2021).
@@ 3 Merton’s portfolio problem: Utility maximisation
$\mathcal{V}_{t}(x,v):=\operatorname*{ess\,sup}_{\alpha\in\mathcal{A}_{t}}\mathbb{E}_{x,v}\!\left[U\!\left(X_{T}^{t,x,v,\alpha}\right)\mid\mathcal{F}_{t}\right],\qquad x>0,v>0,\;t\in[0,T],$ (3.22)
@@
In fact, Fouque and Hu (2019) showed that if the Sharpe-ratio $\lambda$ in 2.14 is bounded and has bounded derivative, then the value process $\mathcal{V}_{t}$ can be expressed as $\mathcal{V}_{t}(x,v)=J_{t}(X_{t}^{\alpha}=x,V_{t}=v)$, where
@@
for the power utility case ($F(t,X_{t}^{\alpha})=U(X_{t}^{\alpha})=\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}$) even if the volatility process $V_{t}$ is non-Markovian. This approach is called the martingale distortion transformation and was first introduced in the seminal paper Zariphopoulou (2001) and later transferred to a non-Markovian setting with Hölder-type inequalities in Tehranchi (2004) for both power utility and exponential utility case (in which $F(t,X_{t}^{\alpha})=-\frac{1}{\gamma}\exp\Big(-\gamma e^{\int^{T}_{t}r(s)ds}X_{t}^{\alpha}\Big)$). The extension to the multi-asset case is straight forward in the case of a bounded risk premium and degenerate correlation structure(cf. Fouque and Hu (2019), Remark 2.5.).
@@
Note that, in contrast to Han and Wong (2020); Aichinger and Desmettre (2021), this representation is expressed in terms of the adjusted forward process, as defined in (2.19)–(2.20), rather than the forward process used in these works. We offer explicit solutions to the optimal portfolio policies that depend on the function $\psi$, solution of the above-mentionned time-dependent multivariate Riccati-Volterra equation. Let $\Lambda$ be defined as
@@
$a(p)\left(|\lambda_{s}|^{2}+\left|\Lambda_{s}\right|^{2}\right)\;=\;a(p)\sum_{i=1}^{d}V_{s}^{i}\left(\theta_{i}^{2}+\nu_{i}^{2}\varsigma^{i}(s)^{2}\psi^{i}(T-s)^{2}\right)\;\leq\;a\sum_{i=1}^{d}V_{s}^{i},$ (3.28)
@@
Power and exponential utility functions are widely adopted in the literature and display distinct risk-aversion properties. Specifically, the power utility function exhibits constant relative risk aversion (CRRA), whereas the exponential utility function is characterized by constant absolute risk aversion (CARA). Optimal strategies corresponding to these utility functions are presented in the sequel, within the framework of the model introduced in Section 2.

### 06 Apr 2026 v2

@@ 1 Introduction
The modeling of asset price dynamics has undergone a paradigm shift with the empirical observation that both implied and realized volatilities of major financial indices exhibit significantly rougher sample paths (Gatheral et al., 2018) than those generated by classical Brownian-motion-based models. This observation has sparked a rapidly expanding body of research on rough volatility models. Motivated by the widespread practical success of the celebrated Heston (1993) stochastic volatility framework, several rough extensions have been proposed, most notably the rough Heston model (El Euch and Rosenbaum, 2019) builds on market microstructure and its generalisation, the Volterra Heston model in Abi Jaber et al. (2019). Recent remarkable advances include the introduction of the so-called fake stationary Volterra Heston models in Gnabeyeu et al. (2025, 2026) with the aim of providing a unified and consistent framework that captures both short- and long-maturity behaviors and enables robust fitting across the entire term structure. This broader class of volatility models, encompassing the aforementioned specifications, is obtained by modeling the volatility process as a stochastic Volterra equation of convolution type with a time-dependent diffusion coefficient. Therefore, this paper focuses on the financial market within the fake stationary affine Volterra model.
@@
Motivated by several important empirical stylized facts about real financial markets such as choice among multiple assets, rough volatility behavior, correlations across stocks or assets and leverage effects (i.e., correlation between a stock and its volatility), multivariate rough volatility models have recently been developed ; see, e.g., Abi Jaber et al. (2019); Tomas and Rosenbaum (2021). In Aichinger and Desmettre (2021), the authors analyze the Merton portfolio problems with power utility for a class of multivariate Volterra Heston models that features both inter-asset correlations and correlation between a stock and its volatility. In the present paper, we solve the Merton portfolio problem for investors with power, exponential and logarithmic utility functions, within the class of the so-called fake stationary multivariate affine Volterra models.
Main contributions. Building upon recent developments in Volterra models (Gnabeyeu et al., 2025, 2026) and motivated by recent works and advances on multivariate Volterra volatility modeling Abi Jaber et al. (2019); Tomas and Rosenbaum (2021); Aichinger and Desmettre (2021), the primary objective of this paper is to advance the literature on utility maximization along two main directions:
@@
Organization of the Work. The outline of the paper is as follows: Section 2 gives an overview of the model which is needed throughout the paper: We introduce the multi asset financial market, where volatility is modeled by a multivariate class of fake stationary Volterra square root process, and we state the optimization problem. For such a market model we consider in section 3 two different approaches to solve the Merton portfolio problem. More precisely, in section 3 we investigate the classical problem of maximizing the expected utility of terminal wealth in a multi-asset fake stationary Volterra–Heston volatility market, for both power, exponential and logarithmic utility preferences. We first adapt the martingale distortion transformation used in Fouque and Hu (2019); Han and Wong (2020) to the (degenerate) multivariate case. However, as it is pointed out in Abi Jaber et al. (2021); Aichinger and Desmettre (2021), this only works if the correlation structure is highly degenerate. Inspired by the techniques used in Bäuerle and Li (2013); Aichinger and Desmettre (2021), we then provide a solution for the Merton portfolio optimization problem for a more general correlation structure using a verification argument. In Section 4, we demonstrate the practical implications of our findings through numerical experiments based on a two-dimensional fake stationary rough Heston model. Finally, Section 5 is devoted to the proofs of the main results.
@@ 2 Preliminaries: Multivariate fake stationary affine Volterra models
$\|b(t,x)\|+\||\sigma(t,x)|\|\leq C_{b,\sigma}(1+\|x\|),\quad\text{for all }t\in[0,T]\text{ and }x\in\mathbb{R}^{d}.$
@@
Assume that $K$ is diagonal with scalar kernels $K_{i}$ on the diagonal for $i=1,\ldots,d$ that is completely monotone on $(0,\infty)$ and satisfies for any $T>0$:
@@
    The kernel $K_{i}$ is strictly positive and fulfills:
@@
Remark: For $i=1,\ldots,d$ , as $K_{i}$ is completely monotone on $(0,\infty)$ and not identically zero, we have that $K_{i}$ is nonnegative, not identically zero, non-increasing and continuous on $(0,\infty)$, and it follows from (B. Gripenberg and Saavalainen, 1990, Theorem 5.5.4) that $K_{i}$ has a resolvent of the first kind $r_{i}$ which is nonnegative and non-increasing in the sense that $s\mapsto r_{i}([s,s+t])$ is non-increasing $\forall t\geq 0$.
In the case of $\alpha$- fractional kernel (corresponding to $K_{i}=K_{\alpha_{i}}$ with $\alpha_{i}\in[\frac{1}{2},1)$), by Gnabeyeu et al. (2026), Equation (2.2) admits at least a unique-in-law positive weak solution as a scaling limit of a sequence of time-modulated Hawkes processes with heavy-tailed kernels in a nearly unstable regime. Moreover, under assumption 2.1 for some $p>0$, a solution $t\mapsto V_{t}^{i}$ to Equation (2.2) starting from $V_{0}^{i}$ has a $\big(\delta_{i}\wedge\theta_{i}\wedge\widehat{\theta}_{i}-\eta\big)$-Hölder pathwise continuous modification on $\mathbb{R}_{+}$ for sufficiently small $\eta>0$ and satisfying (among other properties),
@@
This covers, for instance, constant non-negative kernels, fractional kernels of the form $\frac{t^{\alpha-1}}{\Gamma(\alpha)}\mathbf{1}_{\mathbb{R}_{+}}$ with $\alpha\in(\frac{1}{2},1]$, exponentially decaying kernels ${\rm e}^{-\beta t}$ with $\beta>0$ and more generally the gamma kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}e^{-\beta t}\mathbf{1}_{\mathbb{R}_{+}}$ with $\alpha\in\left(\tfrac{1}{2},1\right]$ and $\beta\geq 0$ ( see e.g. (Gnabeyeu and Pagès, 2025, Propositions 6.1 and 6.3), (Gnabeyeu and Pagès, 2026, Example 2.2 )). These kernels satisfy conditions (2.3)–(2.4), that is, $(\widehat{\cal K}^{cont}_{\widehat{\theta}})$ and $({\cal K}^{cont}_{\theta})$, for $\alpha>1/2$ with $\theta=\widehat{\theta}=\min\bigl(\alpha-\frac{1}{2},\;1\bigr)$.
@@
$\ \mathbb{E}[V_{0}^{i}]=\frac{1-a_{i}}{1-a_{i}\varphi_{\infty}^{i}}\frac{\mu_{\infty}^{i}}{\lambda_{i}}:=v_{\infty}^{i}\qquad\text{and}\qquad\forall\,t\geq 0,\qquad\varphi^{i}(t)=1-\lambda_{i}\int_{0}^{t}K_{i}(t-s)\left(\frac{\mu^{i}(s)}{\lambda_{i}v_{\infty}^{i}}-1\right)\,\,\mathrm{d}s.$ (2.10)
@@
$\textit{($E_{\lambda_{i},c_{i}}$)}:\;\forall\,t\geq 0,\;c_{i}\lambda_{i}^{2}\big(1-(\varphi^{i}(t)-(f_{\lambda_{i}}*\varphi^{i})_{t})^{2}\big)=(f_{\lambda_{i}}^{2}*\varsigma^{i2})(t)\;\textit{where}\;c_{i}=\frac{v_{0}^{i}}{\nu_{i}^{2}v_{\infty}^{i}}\;\textit{i.e.}\;\varsigma^{i}=\varsigma^{i}_{\lambda_{i},c_{i}}.$ (2.12)
Proof : This is a straightforward extension to the multi-dimensional setting of (Gnabeyeu and Pagès, 2025, Proposition 3.4 and Theorem 3.5) (see also (Gnabeyeu et al., 2025, Proposition 4.2 and 4.4)).
@@
Set $E_{D,c}=\bigcup_{i=1}^{d}E_{\lambda_{i},c_{i}}$. From now on, we will assume that there exists a unique positive bounded Borel solution $\varsigma=\varsigma_{D,c}$ on $(0,+\infty)$ of the system of equation $(E_{D,c})$ so that, the corresponding time-inhomogeneous Volterra square root equation (2.2) is refered to as a Multivariate Stabilized Volterra Cox-Ingersoll-Ross (CIR) equation or as a Multivariate fake stationary Volterra CIR equation if, in addition, equation (2.10) holds. The function $\varsigma$ can be interpreted as a control acting on the volatility process (2.2), thereby ensuring that its second moment remains constant over time (see, e.g., Figure 2).
@@ 2.2 Formulation of the stochastic Market model
We assume that $\sigma$ in (2.14) is given by $\sigma=\sigma(V)=\sqrt{\operatorname{diag}(V)}$, where the $\mathbb{R}^{d}_{+}$–valued scaled process $V$ is defined in (2.2) with $\varsigma=\varsigma_{D,c}$ and Equation (2.10) holds true. We will be chiefly interested in the case where $\lambda_{t}$ is linear in $\sigma_{t}$. More specifically, the the market price of risk (risk premium) is assumed to be in the form $\lambda$ == $\big(\theta_{1}\sqrt{V^{1}},\ldots,\theta_{d}\sqrt{V^{d}}\big)^{\top}$, for some constant $\theta_{i}\geq 0$, so that the dynamics for the stock prices (2.14) reads following Kraft (2005); Abi Jaber et al. (2019)
@@
We state the following existence and uniqueness result from Gnabeyeu et al. (2026) which is extended to the multi-dimensional setting.
@@
((Gnabeyeu et al., 2026, Theorem 3.1 and Remark on Theorem 3.2)). Under Assumption 2.1, the stochastic Volterra equation (2.16)-(2.2) has a unique in law continuous $\mathbb{R}^{d}_{+}\times\mathbb{R}^{d}_{+}$-valued weak solution $(S,V)$ for any initial condition $(S_{0},V_{0})\in\mathbb{R}^{d}_{+}\times\mathbb{R}^{d}_{+}$ defined on some filtered probability space $(\Omega,\mathcal{F},(\mathcal{F})_{t\geq 0},\mathbb{P})$ such that
@@
From now on, we set $g_{0}(t):=\varphi(t)V_{0}+\int_{0}^{t}K(t-s)\mu(s)ds$ and $Z_{t}:=\int_{0}^{t}DV_{s}ds+\nu\varsigma(s)\sqrt{\operatorname{diag}(V_{s})}dW_{s},$ for all $t\geq 0$ so that Equation (2.2) reads
@@
This also provides an extension to the inhomogeneous setting of the models considered in Abi Jaber et al. (2019, 2021); Aichinger and Desmettre (2021). Consequently, our main results (Theorems 3.7 and 3.12) extends (Han and Wong, 2020, Theorem 3.3 and Theorem 3.6), (Aichinger and Desmettre, 2021, Theorem 3.2) to the multivariate, time-dependent diffusion coefficient case.
@@ 3 Merton’s portfolio problem: Utility maximisation
$\mathcal{V}_{t}(x,v):=\operatorname*{ess\,sup}_{\alpha\in\mathcal{A}_{t}}\mathbb{E}_{x,v}\!\Big[U\!\Big(X_{T}^{t,x,v,\alpha}\Big)\mid\mathcal{F}_{t}\Big],\qquad x>0,v>0,\;t\in[0,T].$ (3.22)
@@
In fact, Fouque and Hu (2019) showed that if the Sharpe-ratio $\lambda$ in (2.14) is bounded and has bounded derivative, then the value process $\mathcal{V}_{t}$ can be expressed as $\mathcal{V}_{t}(x,v)=J_{t}(X_{t}^{\alpha}=x,V_{t}=v)$, where
@@
for the power utility case (in which $F(t,X_{t}^{\alpha})=U(X_{t}^{\alpha})=\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}$) even if the volatility process $V_{t}$ is non-Markovian. This approach is called the martingale distortion transformation and was first introduced in the seminal paper of Zariphopoulou (2001) and later transferred to a non-Markovian setting with Hölder-type inequalities in Tehranchi (2004) for both power utility and exponential utility case (in which $F(t,X_{t}^{\alpha})=-\frac{1}{\gamma}\exp\Big(-\gamma e^{\int^{T}_{t}r(s)ds}X_{t}^{\alpha}\Big)$). The extension to the multi-asset case is straight forward in the case of a bounded risk premium and degenerate correlation structure(cf. Fouque and Hu (2019), Remark 2.5.).
@@
Note.
It is worth noting that, in contrast to existing literature on utility maximization under Volterra stochastic environments (see, e.g., Han and Wong (2020); Aichinger and Desmettre (2021)), our approach presents the stochastic factor, or auxiliary process $\Gamma$, as the solution to a BSDE. Furthermore, its exponential representation mentioned earlier is expressed in terms of the adjusted forward process, as defined in equations (2.19)–(2.20), rather than the forward process used in these works.
We will offer explicit solutions to the optimal portfolio policies that depend on the function $\psi$, solution of the above-mentionned time-dependent multivariate Riccati-Volterra equation. Let $\Lambda$ be defined as
@@
$a(p)\left(|\lambda_{s}|^{2}+\left|\Lambda_{s}\right|^{2}\right)\;=\;a(p)\sum_{i=1}^{d}V_{s}^{i}\left(\theta_{i}^{2}+\nu_{i}^{2}\varsigma^{i}(s)^{2}\psi^{i}(T-s)^{2}\right)\;\leq\;a\sum_{i=1}^{d}V_{s}^{i},\vskip-2.84544pt$ (3.28)
@@
Power, exponential and logarithmic utility functions are widely adopted in the literature and display distinct risk-aversion properties. Specifically, the power utility function exhibits constant relative risk aversion (CRRA), whereas the exponential utility function is characterized by constant absolute risk aversion (CARA). Optimal strategies corresponding to these utility functions are presented in the sequel, within the framework of the model introduced in Section 2.